OpenStudy (anonymous):

Please HELP!! PROMISE A MEDAL!!! A side of a house is in the shape of a triangle on top of a rectangle. The rectangle is four times as long as it is high, and the altitude of the triangular part is 1 m greater than the height of the rectangle. The total area of the side of the house is 60 square meters. Find the height of this side of the house.

OpenStudy (danjs):

did you draw this thing up and try to put the given information on it yet?

OpenStudy (anonymous):

of course but it doesn't make sense @DanJS

OpenStudy (danjs):

ok let e try and sketch it up here

OpenStudy (danjs):



OpenStudy (danjs):

Let the height of the rectangle be h, the other two measurements relate to that height

OpenStudy (danjs):

It says: The rectangle is four times as long as it is high: h = height we defined length is 4 times that or 4*h

OpenStudy (danjs):

the altitude of the triangular part is 1 m greater than the height of the rectangle height we defined as h so the altitude of the triangle is h + 1

OpenStudy (danjs):

good on how all those measurments were added to the picture?

OpenStudy (anonymous):

yes THANK YOU SO MUCH i feel so stupid now thank you again!!!

OpenStudy (danjs):

What is the formula for the total area of both those figures added together?

OpenStudy (danjs):

Triangle = 1/2 * (base)*height = (1/2)*4h *(h+1) Rectangle = base * height = 4h * h Total Area = Triangle Area + Rectangle Area \[60 = \frac{ 1 }{2 }*4h*(h+1) + 4h*h\]

OpenStudy (danjs):

is that what you did?

OpenStudy (anonymous):

yes thank you again i got the problem

OpenStudy (danjs):

what is h= i just figured it, you get two answers, but only one is reasonable

OpenStudy (anonymous):

h=3

OpenStudy (danjs):

yep same here, you can add those to the diagram as a double check if you want

OpenStudy (danjs):



OpenStudy (danjs):

60 = 12*3 + 0.5*12*4 TRUE, h=3 is right

OpenStudy (danjs):

tag me another if you have more, these are kinda fun probs

OpenStudy (anonymous):

ok i will thank you again